# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ Abstract.
We introduce a new concept named shift system. This is a purely Lie algebraic setting to develop the geometric representation theory of Feigin-Tipunin construction [FT10] of logarithmic conformal field theories. After reformulating the discussion in [Sug21, Sug22] under this new setting, as an application, we extend almost all the main results of these papers to the (multiplet) principal W-algebra at positive integer level associated with a simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{osp}(1|2r)$, respectively.
@@ 1. Introduction
The study of logarithmic conformal field theory (LCFT) has become increasingly important in recent years. However, LCFT is much more complicated than rational cases and has not been well studied. In 2010, Feigin-Tipunin [FT10] proposed a new geometric construction of LCFTs called the Feigin-Tipunin (FT) construction, i.e., the 00-th sheaf cohomology $H^{0}(G\times_{B}V)$ of a $G$-equivariant VOA-bundle $G\times_{B}V$ over the flag variety $G/B$, and claimed that multiplet $W$-algebras  [^1] In the case of $\mathfrak{g}=\mathfrak{sl}_{2}$, this is one of the most famous and well studied LCFTs called triplet Virasoro algebra. Here, we use the term multiplet W-(super)algebra as a generic term for variations of the triplet Virasoro algebra (e.g., the cases of singlet/doublet, higher rank, general W-algebra, etc.). defined by the joint-kernel of screening operators can be constructed and studied by their method. The second author [Sug21, Sug22] gave rigorous proofs of the claims in [FT10] and several fundamental properties on the multiplet $W$-algebra and the corresponding principal $W$-algebra for simply-laced cases. In [CNS24], he also studied the $V^{(p)}$-algebra ($=$ doublet affine $\mathfrak{sl}_{2}$) [Ada16, ACGY21] by combining the method in [Sug21, Sug22] with the inverse quantum Hamiltonian reduction [Ada19]. On the other hand, a close examination of [Sug21, Sug22] implies that, in fact, the VOA-structure does not play an important role, and thus many arguments can be described and developed under a purely Lie algebraic setting. Extracting such Lie algebraic setting and discussion will not only be useful for future studies of LCFTs, but also for people with more general backgrounds, such as representation theory, topology, number theory and physics [GPPV20, CCFGH18, CCKPS24, CFGHP22, Sug].
@@
In Part I of the present paper, we will introduce such a purely Lie algebraic setting, named shift system, and prove that all the main results of [Sug21] can be understood and derived from the setting (see Theorem 1.1). In other words, if one can prove that a VOA-module $V$ (e.g., irreducible lattice VOA-module) fits into a shift system, all the main results of [Sug21] can be used to the corresponding multiplet $W$-algebra. In particular, it has the FT construction $H^{0}(G\times_{B}V)$ and we can use a geometric representation theory. As demonstrated in Section 2.4, the verification of the axiom of shift system is basically reduced to a relatively easy computation ($=$ Serre relation of long screening operators) on $V$ and the rank $1$ case ($=$ “Felder complex” of short screening operators). Hence, the shift system provides a useful framework for reducing the study of higher rank LCFTs to the rank $1$ cases and free field algebras. Moreover, with some information about the corresponding “W-algebra” $H^{0}(G\times_{B}V)^{G}$ (e.g., Kazhdan-Lusztig decomposition), we can also prove all the main results of [Sug22]. Using these results, in Part II, we will give the shift system and the FT constructions corresponding to the (multiplet) principal $W$-(super)algebras $\mathbf{W}^{k}(\mathfrak{g})$ for any simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{g}=\mathfrak{osp}(1|2n)$, respectively, and prove almost all the main results of [Sug21, Sug22] for these new cases. These results in themselves are merely a sequential and relatively straightforward extension of [Sug21, Sug22], and the arguments and methods used are also essentially the same as in these works. However, as mentioned briefly at the end of the previous paragraph, the value of this paper lies in the change of perspective, which greatly increases the transparency and flexibility of the discussion in LCFTs, by focusing not on individual examples of LCFT and their complicated VOA-structures, but on a simple Lie algebraic pattern (i.e., shift system) behind them in common. In that sense, the application to the two examples in Part II is merely a demonstration of this new perspective which will be further developed itself and applied to more LCFTs in the future (see the last of introduction below).
@@
Let us describe the overview of the present paper. For a finite dimensional simple Lie algebra $\mathfrak{g}$, we consider a triple $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, named shift system (see Definition 1.6). Here $\Lambda$ is a module over the Weyl group $W$ of $\mathfrak{g}$, $\uparrow\colon W\times\Lambda\rightarrow\mathfrak{h}^{\ast}$ is a map (shift map) satisfying a few simple axioms, and $V_{\lambda}$ is a graded weight $B$-module with a parameter $\lambda\in\Lambda$ and a “Felder complex” [Fel89] for each direction $i\in\Pi$ (where $\Pi$ is the set of simple roots). Recall that in [FT10, Sug21, Sug22], $V_{\lambda}$ is an irreducible module $V_{\sqrt{p}(Q+\lambda)}$ over the lattice VOA $V_{\sqrt{p}Q}$ with a conformal grading $V_{\lambda}=\bigoplus_{\Delta}V_{\lambda,\Delta}$ and a parameter $\lambda\in\Lambda\simeq(\sqrt{p}Q)^{\ast}/\sqrt{p}Q\simeq\tfrac{1}{p}Q^{\ast}/Q$, and the long/short screening operators define $B$- and $W$-actions on $V_{\lambda}$ and $\Lambda$, respectively. The shift of Cartan weights by the short screening operators satisfies the axioms of the shift map above. The situation in [CNS24] is a little more complicated, but essentially the same. Therefore, the shift system can be regarded as an abstraction of the relationship between such VOA-modules and screening operators acting on them, forgetting the VOA-structure. The first main theorem of Part I asserts that all the main results of [Sug21] hold even under such abstraction.
@@
    (Feigin-Tipunin conjecture/construction) The evaluation map
    $\operatorname{ev}\colon H^{0}(G\times_{B}V_{\lambda})\rightarrow\bigcap_{i\in\Pi}\ker Q_{i,\lambda}|_{V_{\lambda}},\ \ s\mapsto s(\operatorname{id}_{G/B})$
    is injective, and is isomorphic iff $\lambda\in\Lambda$ satisfies the following condition:
    For any $(i,j)\in\Pi\times\Pi$, $\lambda\in\Lambda^{\sigma_{j}}$ or $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=-\delta_{ij}$. (weak)
@@
    (Borel-Weil-Bott type theorem) For a minimal expression $w_{0}=\sigma_{i_{l(w_{0})}}\cdots\sigma_{i_{1}}\sigma_{i_{0}}$ of the longest element $w_{0}$ of $W$ (where $\sigma_{i_{0}}=\operatorname{id}$ for convenience), if $\lambda\in\Lambda$ satisfies the following condition:
    For any $10\leq m\leq N-$ 1, $(\sigma_{i_{m}}\cdots\sigma_{i_{0}}\uparrow\lambda,\alpha_{i_{m+1}}^{\vee})=0$, (strong)
@@
Since the constructions and homomorphisms in Theorem 1.1 are natural, they are compatible with additional algebraic (e.g, VOA-module) structures. In particular, if $V_{0}$ is a VOA and $V_{\lambda}$ are $V_{0}$-modules, then $H^{0}(G\times_{B}V_{0})$ has the induced VOA-structure and $H^{n}(G\times_{B}V_{\lambda})$ are $H^{0}(G\times_{B}V_{0})$-modules (see [Sug22, Section 2.1, Corollary 2.21]). It allows us to develop a geometric representation theory of multiplet $W$-(super)algebras, traditionally defined in the form of the right-hand side of Theorem 1.1(1). Except for rank 1 cases, these VOAs are difficult to study algebraically due to their complicated VOA structures. However, the geometric construction allows us to avoid such difficulties and obtain a variety of results qualitatively.
In Part II, we demonstrate use and utility of the shift system and Theorem 1.1 by applying them to specific VOSAs. Let us consider the rescaled root lattice $\sqrt{p}Q$ for some $p\in\mathbb{Z}_{\geq 1}$. We consider the case where $\sqrt{p}Q$ is positive-definite integral, namely, $p\in r^{\vee}\mathbb{Z}_{\geq 1}$ for the lacing number $r^{\vee}$ or $p$ is odd and $\mathfrak{g}=B_{r}$. In the first case, $\sqrt{p}Q$ is even and we consider $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}$. On the other hand, in the second case, $\sqrt{p}Q$ is odd and we need a modification $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F$ by the free fermion $F$. In Section 2.4, from minimal natural assumptions (2.8) or (2.9), shift systems $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ are uniquely constructed (Theorem 2.13). As noted in the second paragraph above, only relatively easy computations on $V_{\lambda}$ and some basic results in rank $1$ cases ([AM08] and [AM09, AM208]) are used here. Therefore, Theorem 1.1 is applied to our cases (for more detail, see Section 3).
@@
$H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\beta\in P_{+}}L_{\beta}\otimes\mathcal{W}_{-\beta+\lambda},$
where $L_{\beta}$ is the irreducible $\mathfrak{g}$-module with highest weight $\beta$, and $\mathcal{W}_{-\beta+\lambda}$ is the multiplicity of a weight vector of $L_{\beta}$. In Part II, we consider additional VOA structures, where $\mathcal{W}_{0}\simeq H^{0}(G\times_{B}V_{0})^{G}$ is a subalgebra of the multiplet $W$-(super)algebra $H^{0}(G\times_{B}V_{0})$ and $\mathcal{W}_{-\beta+\lambda}$ is a $\mathcal{W}_{0}$-modules. If $\lambda$ satisfies (strong), then by combining $H^{n>0}(G\times_{B}V_{\lambda})\simeq 0$ with the Atiyah-Bott localization formula, we obtain the Weyl-type character formula
$\operatorname{ch}_{q}H^{0}(G\times_{B}V_{\lambda})=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\operatorname{ch}_{q}\mathcal{W}_{-\beta+\lambda}=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\beta-\sigma\uparrow\lambda}.$ (1.1)
On the other hand, by combining the case $n=0$ with Serre duality, we have $H^{0}(G\times_{B}V_{\lambda})\simeq H^{0}(G\times_{B}V_{\lambda^{\prime}})^{\ast}$ for some $\lambda^{\prime}\in\Lambda$. In Part II, it leads the self-duality and simplicity of the vacuum case $H^{0}(G\times_{B}V_{0})$. By the quantum Galois theory [DM97, McR20], $\mathcal{W}_{-\beta}$ are also simple as $\mathcal{W}_{0}$-modules. On the other hand, if a Kazhdan-Lusztig-type character formula for the simple quotient of $\mathcal{W}_{-\beta+\lambda}$ is known, by comparing it with the Weyl-type formula above, we can extend the simplicity to the whole $\lambda\in\Lambda$ satisfying (strong). For more detail, see Section 2.2.
Let us go back to our special cases above. In our cases, $\mathcal{W}_{0}$ is the principal $W$-(super)algebra $\mathbf{W}^{k}(\mathfrak{g})$ or $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))$, and thus we have the Kazhdan-Lusztig type character formula derived from [KT00, KT98] and the exactness of $+$-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ in [Ar07] (however, in the second case we assume the latter). By applying the discussion in the last paragraph, we obtain the following (for more detail, see Theorem 4.3 and 4.7).
@@
(For the case $\mathfrak{g}=ADE$, see [Sug22]) Let us consider the setup in Theorem 2.13. Then for each case, we have $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})$ and $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$, respectively. Furthermore, in the first case, for any $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$, $\mathcal{W}_{-\alpha+\lambda}$ and $H^{0}(G\times_{B}V_{\lambda})$ are simple as $\mathcal{W}_{0}$- and $H^{0}(G\times_{B}V_{0})$-modules, respectively. In the second case, under similar conditions with the restriction $\lambda^{\bullet}=0$ and the assumption that the $+$-reduction is exact, the same simplicity theorem holds.
Finally, let us discuss two directions of future works to which Theorem 1.1 points. The geometric representation theory of Feigin-Tipunin construction has similar aspects to the modular representation theory. Indeed, in our special cases, the strong condition in Lemma 2.9(1) also appear in the Borel-Weil-Bott theorem for ${}^{L}G$ [Jan03, II.5]. Ultimately, such a similarity should come from the log Kazhdan-Lusztig correspondence (e.g., [FGST06, FT10, NT11, MY20, GN24, CLR23]), i.e., an expected categorical equivalence between a multiplet $W$-algebra and a “quantum group”. For example, in our case, $H^{0}(G\times_{B}V_{\sqrt{p}Q})$ is expected to correspond to the small quantum group $u_{\zeta}({}^{L}\mathfrak{g})$ at $\zeta=\frac{\pi i}{p}$.  [^3] On the other hand, our super case is expected to correspond to some small quantum group at $\zeta=\tfrac{2\pi i}{2m-1}$ (see [AM09]). Although it has already been proved [GN24] in the case $\mathfrak{g}=\mathfrak{sl}_{2}$, it seems necessary to develop the geometric method rather than algebraic methods for higher rank generalizations.
On the other hand, under the correspondence of the Borel-Weil-Bott theorems above, Theorem 1.1(1) clearly does not hold in modular representation theory. In this regard, the second author conceived the following idea a few years ago: Let $\vec{\lambda}=(\lambda_{1},\dots,\lambda_{N})\in\Lambda_{1}\times\cdots\times\Lambda_{N}$ be a parameter sequence and $\tilde{V}_{\vec{\lambda}}$ a shift system with respect to $\lambda_{1}\in\Lambda_{1}$. If we consider the situation such that for each $10\leq n\leq N-$ 1, an appropriate quotient
@@
is a shift system with respect to $\lambda_{n+1}\in\Lambda_{n+1}$, then we finally obtain the nested Feigin-Tipunin construction
@@
Because a shift system appears at each stage of nesting, Theorem 1.1 can be applied repeatedly. In particular, repeated application of (1.1) 444 To compute $\operatorname{ch}_{q}\tilde{H}^{0}(G\times_{B}\cdots)$ from $\operatorname{ch}_{q}H^{0}(G\times_{B}\cdots)$ using (1.1), we assume a suitable match between them. in the case $\mathfrak{g}=\mathfrak{sl}_{2}$, surprisingly, almost coincides with the $\hat{Z}$-invariant (or homological block) [GPPV20] of a certain $3$-manifold. This suggests that such nested FT constructions yield rich examples of LCFT whose representation theories are to some extent reducible and controllable by the theory of shift system, providing a new approach for the studies in [CCFGH18, CCKPS24, CFGHP22], etc. In [Sug], the second author will give a combinatorial algorithm that achieves the nested FT construction in the case $\mathfrak{g}=\mathfrak{sl}_{2}$.
Acknowledgments H.L. thanks to Antun Milas for motivating questions, and to Drazen Adamovic and Naoki Genra for useful discussions. S.S. thanks to Naoki Genra and Hiroshi Yamauchi for useful comments on naming. The authors thanks to Tomoyuki Arakawa for useful feedback. The majority of this work was done when the authors were postdocs in YMSC, Tsinghua University, and they are supported by Beijing Natural Science Foundation grant IS24011. S.S. is also supported by Caltech-Tsinghua joint Postdoc Fellowships.
@@
In Part I, we describe the discussion of [FT10, Sug21, Sug22] in an axiomatic (in other words, purely Lie algebraic) manner. We also generalize the discussion in these papers to include non-simply laced cases.

### v3

@@ Abstract.
We introduce a purely Lie algebraic formalization of the Feigin–Tipunin’s geometric construction of logarithmic CFTs/VOAs. After reformulating the geometric representation theory of FT construction under this new setting, within this framework, we uniformly construct the (multiplet) principal W-algebras at positive integer level associated with any simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{osp}(1|2r)$, thereby establishing Weyl-type character formulas and simplicity theorems that extend the second author’s previous results.
@@ 1. Introduction
The study of logarithmic conformal field theory (LCFT) has become increasingly important in recent years. However, LCFT is much more complicated than rational cases and has not been well studied. In 2010, Feigin–Tipunin [27] proposed a geometric construction of LCFTs called the Feigin–Tipunin (FT) construction, i.e., the 0-th sheaf cohomology $H^{0}(G\times_{B}V)$ of a $G$-equivariant VOA-bundle $G\times_{B}V$ over the flag variety $G/B$, and claimed that multiplet W-algebras [^1] In the case of $\mathfrak{g}=\mathfrak{sl}_{2}$ with the principal nilpotent element, this is one of the most famous and well studied LCFTs called triplet Virasoro algebra (see e.g. [7]). Here the “multiplet W-(super)algebra” is intended to be a generic term for variations of the triplet Virasoro algebra (e.g., the cases of singlet/doublet, higher rank, general W-algebra, etc.). defined by the joint-kernel of screening operators can be constructed and studied by their method. The second author [45, 46] gave rigorous proofs of the claims in [27] and several fundamental properties on the multiplet W-algebra and the corresponding principal W-algebra for simply-laced cases. In [14], he also studied the $V^{(p)}$-algebra (== doublet affine $\mathfrak{sl}_{2}$) [1, 4] by combining the method in [45, 46] with the inverse quantum Hamiltonian reduction [2]. On the other hand, a close examination of [45, 46] suggests that, although formulated in the framework of VOAs, the essential arguments ultimately reduce to properties of an underlying Lie algebraic setting. As detailed below, extracting such Lie algebraic setting and argument not only provides a clear perspective for the studies of LCFTs, but also sheds light [47, 48] on the expected correspondence between LCFTs and $3$-manifolds via the $\hat{Z}-$invariants [15, 16, 17, 30].
@@
In Part I, we will introduce such a purely Lie algebraic setting, tentatively named shift system, and prove that the main results of [45] can be understood and derived from the setting (Theorem 1.1). In other words, if a VOA-module $V$ (e.g. irreducible lattice VOA-module) fits into a shift system, the corresponding multiplet W-algebra satisfies Theorem 1.1. In particular, as it has the FT construction $H^{0}(G\times_{B}V)$, we can study it using the geometric representation theory. As demonstrated in Section 2.4, the verification of the axiom of shift system is basically reduced to a relatively easy computation (== Serre relation of long screening operators) on $V$ and the rank $1$ case (== “Felder complex” of short screening operators). Hence, the shift system provides a useful framework for reducing the study of higher rank LCFTs to much easier cases: the rank $1$ cases and free field algebras. Moreover, with some information about the corresponding “W-algebra” $H^{0}(G\times_{B}V)^{G}$ (e.g. Kazhdan–Lusztig decomposition), the main results of [46] are proved (Section 2.2). Using these results, in Part II, we will give the shift system and the FT constructions corresponding to the (multiplet) principal W-(super)algebras $\mathbf{W}^{k}(\mathfrak{g})$ for any simple Lie algebra $\mathfrak{g}$ and Lie superalgebra $\mathfrak{g}=\mathfrak{osp}(1|2n)$, respectively, and prove almost all the main results of [45, 46] for these new cases (Therem 1.2). At first glance, these results themselves may appear to be just a straightforward extension of [45, 46]. However, as mentioned briefly at the end of the last paragraph, the value of this paper lies in the change of perspective, which greatly increases the transparency and flexibility of the discussion in LCFTs, by focusing not on individual examples of LCFT and their complicated VOA-structures but on a simple Lie algebraic pattern behind them in common. In that sense, the application to the two examples in Part II is a demonstration of this new perspective which will be further developed itself and applied to much broader class of expected LCFTs corresponding to $3$-manifolds in the future (see the last of introduction below).
@@
Let us describe the overview of this paper. For a finite dimensional simple Lie algebra $\mathfrak{g}$ of rank $r$, we consider a triple $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$, named shift system (Definition 1.6). Here $\Lambda$ is a module over the Weyl group $W$ of $\mathfrak{g}$, $\uparrow\colon W\times\Lambda\rightarrow\mathfrak{h}^{\ast}$ is a map (shift map) satisfying a few simple axioms, and $V_{\lambda}$ is a graded weight $B$-module with a parameter $\lambda\in\Lambda$ and a “Felder complex” [21] for each direction $i\in I:=\{1,\dots,r\}$. Recall that in [27, 45, 46], $V_{\lambda}$ is an irreducible module $V_{\sqrt{p}(Q+\lambda)}$ over the lattice VOA $V_{\sqrt{p}Q}$ with a conformal grading $V_{\lambda}=\bigoplus_{\Delta}V_{\lambda,\Delta}$ and a parameter $\lambda\in\Lambda\simeq(\sqrt{p}Q)^{\ast}/\sqrt{p}Q\simeq\tfrac{1}{p}Q^{\ast}/Q$, and the long/short screening operators define $B$- and $W$-actions on $V_{\lambda}$ and $\Lambda$, respectively. The shift of Cartan weights by the short screening operators satisfies the axioms of the shift map above. The situation in [14] is a little more complicated, but essentially the same. Therefore, the shift system can be regarded as an abstraction of the relationship between such VOA-modules and screening operators acting on them, forgetting the VOA-structure. The first main theorem of Part I asserts that all the main results of [45] hold even under such abstraction.
@@
    (Feigin–Tipunin construction) The evaluation map
    $\operatorname{ev}\colon H^{0}(G\times_{B}V_{\lambda})\rightarrow\bigcap_{i\in I}\ker Q_{i,\lambda}|_{V_{\lambda}},\ \ s\mapsto s(\operatorname{id}_{G/B})$
    is injective, and is isomorphic iff $\lambda\in\Lambda$ satisfies the following “weak condition”:
    (weak) For any $(i,j)\in I\times I$, $\lambda\in\Lambda^{\sigma_{j}}$ or $({\sigma_{j}}\uparrow{\lambda},\alpha_{i}^{\vee})=-\delta_{ij}$.
@@
    (Borel–Weil–Bott-type theorem) For a minimal expression $w_{0}=\sigma_{i_{l(w_{0})}}\cdots\sigma_{i_{1}}\sigma_{i_{0}}$ of the longest element $w_{0}$ of $W$ (where $\sigma_{i_{0}}=\operatorname{id}$ for convenience), if $\lambda\in\Lambda$ satisfies the following “strong condition”:
    (strong) For any $0\leq m\leq N-1$, $(\sigma_{i_{m}}\cdots\sigma_{i_{0}}\uparrow\lambda,\alpha_{i_{m+1}}^{\vee})=0$,
@@
Since the constructions and homomorphisms in Theorem 1.1 are natural, they are compatible with additional algebraic (e.g, VOA-module) structures. In particular, if $V_{0}$ is a VOA and $V_{\lambda}$ are $V_{0}$-modules, then $H^{0}(G\times_{B}V_{0})$ has the induced VOA-structure and $H^{n}(G\times_{B}V_{\lambda})$ are $H^{0}(G\times_{B}V_{0})$-modules (see [46, Section 2.1, Corollary 2.21]). It allows us to develop a geometric representation theory of multiplet W-(super)algebras, traditionally defined in the form of the right-hand side of Theorem 1.1(1). Except for rank 1 cases, these VOAs are difficult to study algebraically due to their complicated VOA structures. However, the geometric construction allows us to avoid the bottleneck and obtain a variety of results qualitatively, as in [45, 46, 14].
In Part II, we demonstrate use and utility of the shift system and Theorem 1.1 by applying them to specific VOSAs. Let us consider the rescaled root lattice $\sqrt{p}Q$ for some $p\in\mathbb{Z}_{\geq 1}$. We consider the case where $\sqrt{p}Q$ is positive-definite integral, namely, $p\in r^{\vee}\mathbb{Z}_{\geq 1}$ for the lacing number $r^{\vee}$ or $p$ is odd and $\mathfrak{g}=B_{r}$. In the first case, $\sqrt{p}Q$ is even and we consider $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}$. On the other hand, in the second case, $\sqrt{p}Q$ is odd and we need a modification $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F$ by the free fermion $F$. In Section 2.4, from minimal natural assumptions (2.8) or (2.9), shift systems $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ are uniquely constructed (Theorem 2.13). As noted in the second paragraph above, only relatively easy computations on $V_{\lambda}$ and some basic results in rank $1$ cases ([7] and [9, 8]) are used here. Therefore, Theorem 1.1 is applied to our cases (for more detail, see Section 3).
@@
(1.1) $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\beta\in P_{+}}L_{\beta}\otimes\mathcal{W}_{-\beta+\lambda},$
where $L_{\beta}$ is the irreducible $\mathfrak{g}$-module with highest weight $\beta$, and $\mathcal{W}_{-\beta+\lambda}$ is the multiplicity of a weight vector of $L_{\beta}$. In Part II, we consider additional VOA structures, where $\mathcal{W}_{0}\simeq H^{0}(G\times_{B}V_{0})^{G}$ is a subalgebra of the multiplet W-(super)algebra $H^{0}(G\times_{B}V_{0})$ and $\mathcal{W}_{-\beta+\lambda}$ is a $\mathcal{W}_{0}$-modules. If $\lambda$ satisfies (strong), then by combining $H^{n>0}(G\times_{B}V_{\lambda})\simeq 0$ with the Atiyah–Bott localization formula, we obtain the Weyl-type character formula
(1.2) $\operatorname{ch}_{q}H^{0}(G\times_{B}V_{\lambda})=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\operatorname{ch}_{q}\mathcal{W}_{-\beta+\lambda}=\sum_{\beta\in P_{+}}\operatorname{dim}L_{\beta}\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\beta-\sigma\uparrow\lambda}.$
On the other hand, by combining the case $n=0$ with Serre duality, we have $H^{0}(G\times_{B}V_{\lambda})\simeq H^{0}(G\times_{B}V_{\lambda^{\prime}})^{\ast}$ for some $\lambda^{\prime}\in\Lambda$. In the VOA side, it leads the self-duality and simplicity of the vacuum case $H^{0}(G\times_{B}V_{0})$. By the quantum Galois theory [19, 43], $\mathcal{W}_{-\beta}$ are also simple as $\mathcal{W}_{0}$-modules. In addition, if a Kazhdan–Lusztig-type character formula for the simple quotient of $\mathcal{W}_{-\beta+\lambda}$ is known, by comparing it with (1.2), we can extend the simplicity to the whole $\lambda\in\Lambda$ satisfying (strong). For more detail, see Section 2.2.
Let us go back to our special cases above. In our cases, $\mathcal{W}_{0}$ is the principal W-(super)algebra $\mathbf{W}^{k}(\mathfrak{g})$ or $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))$, and thus we have the Kazhdan–Lusztig type character formula derived from [37, 38] and the exactness of ++-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ in [10] (however, in the second case we assume the latter). By applying the discussion in the last paragraph, we can extend the main results of [5] to our cases (Theorem 4.3 and 4.8).
@@
Let us consider the setup in Theorem 2.13 and Section 4.1.
Then for each case, we have $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{g})\simeq\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})$ and $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$, respectively. Furthermore, in the first case, for any $\alpha\in P_{+}\cap Q$ and $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$, $\mathcal{W}_{-\alpha+\lambda}$ and $H^{0}(G\times_{B}V_{\lambda})$ are simple as $\mathcal{W}_{0}$- and $H^{0}(G\times_{B}V_{0})$-modules, respectively. In the second case, under similar conditions with the restriction $\lambda^{\bullet}=0$ and the assumption that the ++-reduction is exact, the same simplicity theorem holds. In the first case, the above simplicity of $\mathcal{W}_{-\alpha+\lambda}$ also leads to the simplicity and duality of the Arakawa–Frenkel modules $\check{\mathbf{T}}^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}\simeq\mathbf{T}^{1/p}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}$.
The decomposition (1.1) with Theorem 1.2 is regarded as the Schur-Weyl type duality for the multiplet principal W-algebras. In the non-simply laced case, in light of [43], the tensor category of $\mathbf{W}^{k}(\mathfrak{g})$-modules is expected to exhibit a symmetry governed by the Langlands dual Lie algebra $^{L}\mathfrak{g}$.
Finally, let us discuss two directions to which Theorem 1.1 points. Our geometric representation theory of FT construction has similar aspects to the modular representation theory. Indeed, in our special cases, the strong condition in Lemma 2.9(1) also appear in the BWB theorem for $^{L}G$ [32, II.5]. Ultimately, such a similarity should come from the log Kazhdan–Lusztig correspondence (e.g. [24, 27, 44, 42, 29, 13]): an expected categorical equivalence between a multiplet $W$-algebra and a “quantum group”. For example, in our case, $H^{0}(G\times_{B}V_{\sqrt{p}Q})$ is expected to correspond to the small quantum group $u_{\zeta}({}^{L}\mathfrak{g})$ at $\zeta=\frac{\pi i}{p}$. [^3] On the other hand, our super case is expected to correspond to some small quantum group at $\zeta=\tfrac{2\pi i}{2m-1}$ (see [9]). Although the case $\mathfrak{g}=\mathfrak{sl}_{2}$ was already treated in [29], it seems necessary to develop both our geometric methods and W-algebraic approaches for higher rank generalizations (see Remark 4.4).
On the other hand, under the correspondence of the BWB-type theorems above, Theorem 1.1(1) clearly does not hold in the modular representation theory [32]. In this regard, the second author conceived the following idea a few years ago: Let $\vec{\lambda}=(\lambda_{1},\dots,\lambda_{N})\in\Lambda_{1}\times\cdots\times\Lambda_{N}$ be a parameter sequence and $\tilde{V}_{\vec{\lambda}}$ a shift system with respect to $\lambda_{1}\in\Lambda_{1}$. If we consider the situation such that for each $0\leq n\leq N-1$, an appropriate quotient
@@
is a shift system with respect to $\lambda_{n+1}\in\Lambda_{n+1}$, then we finally obtain the nested FT construction
@@
Because a shift system appears at each stage of nesting, Theorem 1.1 can be applied repeatedly. In particular, repeated application of (1.2)444 To compute $\operatorname{ch}_{q}\tilde{H}^{0}(G\times_{B}\cdots)$ from $\operatorname{ch}_{q}H^{0}(G\times_{B}\cdots)$ using (1.2), we assume a suitable match between them. in the case $\mathfrak{g}=\mathfrak{sl}_{2}$, surprisingly, almost coincides with the $\hat{Z}-$invariant [30] of the negative definite Seifert $3$-manifold! This suggests that such nested FT constructions yield rich examples of LCFT whose representation theories are to some extent reducible and controllable by the geometric representation theory of FT construction, providing a blueprint for the “dictionary” between LCFTs and negative definite plumbed $3$-manifolds expected in [15, 16, 17], etc. For $\mathfrak{g}=\mathfrak{sl}_{2}$, the second author already achieved the nested FT construction and reconstructed the $\hat{Z}-$invariants of these $3$-manifolds [47, 48].
Acknowledgments: H.L. thanks to Antun Milas for motivating questions, and to Drazen Adamovic and Naoki Genra for useful discussions. S.S. thanks to Naoki Genra and Hiroshi Yamauchi for useful comments on naming. The authors thank to Tomoyuki Arakawa for useful feedback. The majority of this work was done when the authors were postdocs in YMSC, Tsinghua University.
@@
In Part I, we describe the discussion of [27, 45, 46] in an axiomatic (in other words, purely Lie algebraic) manner. We also generalize the discussion in these papers to include non-simply laced cases.